OpenStudy (anonymous):

Given that sin(2θ)=2/3, the value of sin^6θ+cos^6θ can be written as a/b with a and b coprime positive integers. Find a+b.

OpenStudy (aravindg):

sin(2θ)=23 :O

OpenStudy (anonymous):

are u sure sin(2 theeta) is 23 ?

OpenStudy (anonymous):

no, actually i have now edited theques

OpenStudy (anonymous):

\[\sin(2\theta)=\frac{2}{3}\] \[\sqrt{1-\cos^{2}2\theta }=\frac{2}{3}\] \[1-\cos^{2}2\theta=\frac{4}{9}\] \[\cos^{2}2\theta=\frac{5}{9}\] \[\cos2\theta=\frac{\sqrt{5}}{3}\] \[1-2\sin^{2}\theta=\frac{\sqrt{5}}{3}\] \[2\sin^{2}\theta=1-\frac{\sqrt{5}}{3}\] \[\sin^{2}\theta=\frac{3-\sqrt{5}}{6}\] \[\sin^{6}\theta +\cos^{6}\theta = \sin^{6}\theta + (\sqrt{1-\sin^{2}\theta})^{6}\] \[\sin^{6}\theta+(1-\sin^{2})^{3}\] \[\sin^{2}\theta+1-\sin^{6}\theta+3(1-\sin^{2}\theta)(-\sin^{2}\theta)\] \[1-3\sin^{2}\theta+3\sin^{4}\theta\] Substituting, \[1-3(\frac{ 3-\sqrt{5} }{ 6 })+3(\frac{3-\sqrt{5}}{6})^{2}\] \[1-(\frac{9-3\sqrt{5}}{6})+3(\frac{9+5-6\sqrt{5}}{36})\] \[\frac{6-9+3\sqrt{5}+7-3\sqrt{5}}{6}\] \[\frac{4}{6}\] \[\frac{2}{3}\] so, a+b= 2+3 =5